OpenStudy (anonymous):

What is the equation of the line that passes through (12,4) and is perpendicular to the graph of y=-(3/4)x=2?

OpenStudy (anonymous):

are you sure that equation of graph you have provided is correct there are two equals to sign

OpenStudy (anonymous):

I am sorry it is y=-(3/4)x-2

OpenStudy (anonymous):

Can I use the equation y – y1 = m(x – x1)

OpenStudy (anonymous):

the slope of equation you have given is -3/4 by comparing with y=mx+c as the line we need to find is perpendicular to line you have given, the slope of line we need to find is 4/3 as m1 . m2 = -1 now we have slope = 4/3 and point =(12,4) now you cna use point slope form which i guess you know

OpenStudy (anonymous):

yes y – y1 = m(x – x1) is the one you should use

OpenStudy (anonymous):

Thanks!